### v1

# Littlewood-Offord problems for Ising models

Abstract
We consider the one dimensional Littlewood-Offord problem for general Ising models. More precisely, consider the concentration function
$Q_{n}(x,v)=P(\sum_{i=1}^{n}\varepsilon_{i}v_{i}\in(x-1,x+1)),$
where $x\in\mathbb{R}$, $v_{1},v_{2},\ldots,v_{n}$ are real numbers such that $|v_{1}|\geq 1,|v_{2}|\geq 1,\ldots,|v_{n}|\geq 1$, and $(\varepsilon_{i})_{i=1,2,\ldots,n}$ are spins of some Ising model. Let $Q_{n}=\sup_{x,v}Q_{n}(x,v)$. Under natural assumptions, we show that there exists a universal constant $C$ such that for all $n\geq 1$,
$\binom{n}{[n/2]}2^{-n}\leq Q_{n}\leq Cn^{-\frac{1}{2}}.$
1 Introduction
The classical Littlewood-Offord problem is about the uniform upper bound of the concentration probability
$P(\sum_{i=1}^{n}\varepsilon_{i}v_{i}\in(x-1,x+1))$
where $x$ is a real number, $v_{1},v_{2},\ldots,v_{n}$ are real numbers such that $|v_{1}|\geq 1,|v_{2}|\geq 1,\ldots,|v_{n}|\geq 1$, and $(\varepsilon_{i})_{i=1,2,\ldots,n}$ are independent Rademacher random variables such that for $i=1,2,\ldots,n$, $P(\varepsilon_{i}=1)=P(\varepsilon_{i}=-1)=1/2$. This problem was first considered by Littlewood and Offord [LO43] with applications in the study of real roots of random polynomials. Later, Erdős [Erd45] obtained the following sharp result:
$\sup_{x\in\mathbb{R}}\sup_{v_{1},v_{2},\ldots,v_{n}\in(-\infty,-1]\cup[1,\infty)}P\left(\sum_{i=1}^{n}\varepsilon_{i}v_{i}\in(x-1,x+1)\right)=\binom{n}{[n/2]}2^{-n}.$
The same upper bound was obtained by Kleitman [Kle70] for vectors $x$ and $(v_{i})_{i=1,2,\ldots,n}$. We refer to the reference in [TV12] for the series of work in high dimensions. The inverse problem was first considered by Tao and Vu [TV09]. The inverse problem is closely related to random matrices.
Beyond Rademacher series, we notice the results [JK21] and [Sin22] on similar problems for i.i.d. Bernoulli random variables. Beyond independence, we are aware of the result [Rao21] for symmetric random variables driven by a finite state reversible stationary Markov chain. Together with Peng, we [CP24] considered the Littlewood-Offord problem for spins of Curie-Weiss models, and obtained some sharp and asymptotic results. In the present paper, we consider the Littlewood-Offord problems for spins of Ising models.
Ising models are models for magnets in statistic physics. Let’s briefly introduce Ising models in the following. Consider a graph $G=(V,\mathcal{E})$. Let $V_{n}\subset V$ be a subset of vertices with $n$ vertices. Let $\mathcal{E}_{n}$ be the set of edges adjacent to some vertex in $V_{n}$, i.e.
$\mathcal{E}_{n}=\{\{i,j\}\in\mathcal{E}:i\in V_{n}\text{ or }j\in V_{n}\}.$
Let $\Sigma_{n}=\{-1,1\}^{V_{n}}$ be the configuration space of spins. Let $\kappa\in\{-1,1\}^{V\setminus V_{n}}$ be a boundary condition. Let $h\in\mathbb{R}^{V_{n}}$ be the external field. For a configuration $\sigma\in\Sigma_{n}$, we extend $\sigma$ to a boolean function on $V$ by taking $\sigma_{j}=\kappa_{j}$ for $j\notin V_{n}$. We define the energy of $\sigma$ via a function called Hamiltonian as follows:
$H_{n,J,h,\kappa}(\sigma)=-\sum_{\{i,j\}\in\mathcal{E}_{n}}J_{ij}1_{\sigma_{i}=\sigma_{j}}-\sum_{i\in V_{n}}h_{i}\sigma_{i},$ (1)
where $J_{ij}=J_{ji}\in\mathbb{R}$ is the coupling constant. Without loss of generality, we always assume that
$J_{ij}\neq 0,\forall\{i,j\}\in\mathcal{E}.$
Definition 1.1.
The Gibbs distribution of the Ising model is defined by
$\mu_{n}(\sigma)=e^{-H_{n,J,h,\kappa}(\sigma)}/Z_{n},$ (2)
where $Z_{n}$ is the normalization constant.
Let $(\varepsilon_{i})_{i=1,2,\ldots,n}$ be some random vector sampled according to the Gibbs distribution $\mu_{n}$. We are interested in
$Q_{n}=\sup_{x\in\mathbb{R}}\sup_{|v_{1}|,|v_{2}|,\ldots,|v_{n}|\geq 1}P(\sum_{i=1}^{n}\varepsilon_{i}v_{i}\in(x-1,x+1)).$ (3)
Our main result is the upper and lower bound for $Q_{n}$ under the boundedness assumption on the coupling constants $(J_{ij})$ and the external field $(h_{i})$.
Theorem 1.1.
Assume that
$K=\sup_{n\geq 1}\max_{i\in V_{n}}\sum_{j\in V}|J_{ij}|+\max_{i\in V_{n}}|h_{i}|<+\infty.$ (4)
Then, there exists a universal constant $C(K)$ such that
$\binom{n}{[n/2]}2^{-n}\leq Q_{n}\leq C(K)n^{-\frac{1}{2}}.$ (5)
So, under the boundedness assumption (4), $Q_{n}$ is of the order $1/\sqrt{n}$.
Remark 1.1.
The lower bound $\binom{n}{[n/2]}2^{-n}\leq Q_{n}$ in (5) is quite general and holds for all random vectors $(\varepsilon_{i})_{i}\in\{-1,1\}^{n}$, see Theorem 3.1.
2 Preliminary: Edwards-Sokal coupling for Ising models
In this section, we will explain the Edwards-Sokal coupling [ES88] for Ising models beyond ferromagnetism. For simplicity of notation, we consider the case without external field $h$. The presence of external field $h$ could be transformed to certain boundary condition. Indeed, let $\widetilde{V}_{n}$ be a copy of $V_{n}$, where the vertex $\widetilde{i}\in\widetilde{V}_{n}$ is a copy of the vertex $i\in V_{n}$. Let $\widetilde{V}=V\cup\widetilde{V}_{n}$ be an enlargement of the original vertex set $V$. We enlarge the edge set $\mathcal{E}$ to $\widetilde{\mathcal{E}}$ by adding the edges $\{i,\widetilde{i}\}$ between $V_{n}$ and $\widetilde{V}_{n}$. We put no edge between vertices of $\widetilde{V}_{n}$. The coupling constant $J_{i\widetilde{i}}$ is set to $2h_{i}$ for $i\in V_{n}$ and $\widetilde{i}\in\widetilde{V}_{n}$. We extend the boundary condition $\kappa$ to $\widetilde{\kappa}$ by setting $\widetilde{\kappa}_{\widetilde{i}}=1$ for all $\widetilde{i}\in\widetilde{V}_{n}$. Then, the original Gibbs distribution is the same as the Gibbs distribution on the extended graph with boundary condition $\widetilde{\kappa}$ and vanishing external field. Thus, without loss of generality, we assume that there is no external field.
Define
$\mathcal{E}_{n}^{+}=\{\{i,j\}\in\mathcal{E}_{n}:J_{ij}>0\},\quad\mathcal{E}_{n}^{-}=\{\{i,j\}\in\mathcal{E}_{n}:J_{ij}<0\},$
i.e. $\mathcal{E}_{n}^{+}$ (resp. $\mathcal{E}_{n}^{-}$) is the collection of edges adjacent to $V_{n}$ with positive (resp. negative) coupling constants. Recall that we assume that $J_{ij}\neq 0$ if $\{i,j\}$ is an edge. Hence, $\mathcal{E}_{n}$ is the disjoint union of $\mathcal{E}_{n}^{+}$ and $\mathcal{E}_{n}^{-}$. We define a model with coupled random spins and random edges as follows: Let $\sigma=(\sigma_{i})_{i}\in\{-1,1\}^{V_{n}}$ be the configuration of spins on vertices and let $e=(e_{ij})_{\{i,j\}}\in\{0,1\}^{\mathcal{E}_{n}}$ be the edge configuration. (Here, the edge $\{i,j\}$ is open if and only if $e_{ij}=1$. The edge $\{i,j\}$ is closed if it is not open.) Define the probability $\mu_{FKSW}(\sigma,e)$ that $\sigma$ and $e$ are realized by
$\mu_{FKSW}(\sigma,e)=$ $Z_{FKSW}^{-1}\prod_{\{i,j\}\in\mathcal{E}_{n}^{+}}((1-p_{ij})1_{e_{ij}=0}+p_{ij}1_{e_{ij}=1}1_{\sigma_{i}=\sigma_{j}})$
×$\times\prod_{\{i,j\}\in\mathcal{E}_{n}^{-}}((1-p_{ij})1_{e_{ij}=0}+p_{ij}1_{e_{ij}=1}1_{\sigma_{i}\neq\sigma_{j}}),$ (6)
where $p_{ij}=1-e^{-|J_{ij}|}$ and $Z_{FKSW}$ is the normalizing constant. Note that $\mu_{FKSW}(\sigma,e)>0$ iff the following constraint is satisfied: for all $\{i,j\}\in\mathcal{E}_{n}^{+}$, $e_{ij}$ must be 00 as long as $\sigma_{i}\neq\sigma_{j}$; and for all $\{i,j\}\in\mathcal{E}_{n}^{-}$, $e_{ij}$ must be 00 as long as $\sigma_{i}=\sigma_{j}$. The marginal distribution of the spin configuration $\sigma$ is given by
$\mu_{\mathrm{Ising}}(\sigma)=Z^{-1}_{\mathrm{Ising}}e^{\sum_{\{i,j\}\in\mathcal{E}_{n}^{+}}J_{ij}(1_{\sigma_{i}=\sigma_{j}}-1)}e^{\sum_{\{i,j\}\in\mathcal{E}_{n}^{-}}J_{ij}1_{\sigma_{i}=\sigma_{j}}},$ (7)
which is proportional to
$e^{\sum_{\{i,j\}\in\mathcal{E}_{n}}J_{ij}1_{\sigma_{i}=\sigma_{j}}}.$
Hence, the marginal distribution $\mu_{\mathrm{Ising}}$ is the same as the Gibbs measure of the Ising model (without external field). The marginal distribution of the edge configuration will be called random cluster model. Due to possible presence of negative coupling constants $J_{ij}$, the description of the random cluster model is more complicated. As we don’t need the precise definition of the random cluster model, we will not discuss the details here.
The two conditional distributions $\mu_{FKSW}(e|\sigma)$ and $\mu_{FKSW}(\sigma|e)$ are important to us. We will discuss them in the following.
Given the spin configuration $\sigma$, the edge random variables $(e_{ij})_{ij}$ are conditionally independent. Moreover, for $\{i,j\}\in\mathcal{E}_{n}^{+}$, when $\sigma_{i}\neq\sigma_{j}$, we must have $e_{ij}=0$; when $\sigma_{i}=\sigma_{j}$, the conditional probability $\mu_{FKSW}(e_{ij}=1|\sigma)=p_{ij}$. For $\{i,j\}\in\mathcal{E}_{n}^{-}$, when $\sigma_{i}=\sigma_{j}$, we must have $e_{ij}=0$; when $\sigma_{i}\neq\sigma_{j}$, the conditional probability $\mu_{FKSW}(e_{ij}=1|\sigma)=p_{ij}$. In summary, given the spin configuration $\sigma$, the process $(e_{ij})_{ij}$ is a Bernoulli bond percolation.
Given an edge configuration $e$ with strictly positive possibility, we first define clusters as connected components of $\overline{V_{n}}$ under certain equivalence relation $\sim$ on $\overline{V_{n}}$, where
$\overline{V_{n}}=\{i\in V:i\in V_{n}\text{ or }\exists j\in V_{n}\text{ such that }\{i,j\}\in\mathcal{E}_{n}\}.$
Hence, $\overline{V_{n}}$ is the set $V_{n}$ plus its neighbor vertices. For two different vertices $i_{0}$ and $i_{m}$ in $\overline{V_{n}}$, we define $i_{0}\sim i_{m}$ iff there exists $i_{1},i_{2},\ldots,i_{m-1}\in\overline{V_{n}}$ such that $e_{i_{0}i_{1}}=e_{i_{1}i_{2}}=\cdots=e_{i_{m-1}i_{m}}=1$. Let $C_{1},C_{2},\ldots,C_{\ell},C_{\ell+1},\ldots,C_{\ell+q}$ be the clusters. Here, without loss of generality, we assume that $C_{1},C_{2},\ldots,C_{\ell}$ are contained in $V_{n}$, and $C_{\ell+1},\ldots,C_{\ell+q}$ are not entirely contained in $V_{n}$. To distinguish these two kinds of clusters, we call the former inner clusters and the later boundary clusters. Fix a total order on $V$. For each $k\geq 1$, choose the smallest vertex $c_{k}$ in the cluster $C_{k}$. Define $\eta_{k}=\sigma_{c_{k}}$ for $k\geq 1$. For each $k$, once the spin $\eta_{k}$ is fixed, the spins of the other vertices in $C_{k}$ are fixed at the same time, according to the following rule: If $\{i,j\}\in\mathcal{E}_{n}^{+}$ and $e_{ij}=1$, then $\sigma_{i}=\sigma_{j}$; if $\{i,j\}\in\mathcal{E}_{n}^{-}$ and $e_{ij}=1$, then $\sigma_{i}=-\sigma_{j}$. By similar reasons, the spins of vertices in boundary clusters are determined by the boundary condition. Note that there is no conflict of spins since the edge configuration occurs with strictly positive probability under $\mu_{FKSW}$. Next, we call $\eta_{k}$ the cluster-spin of $C_{k}$. For a vertex $i$ in a cluster $C_{k}$, we define
$S(i)=\eta_{k}\sigma_{i}.$ (8)
In this way, we define a boolean function $S$ on $\overline{V_{n}}$. Although $S$ depends on $e$ and $\sigma$ by definition, it actually depends only on $e$. Finally, given the edge configuration $e$, the cluster-spins $\eta_{\ell+1},\ldots,\eta_{\ell+q}$ of boundary clusters are determined by the boundary condition, and the cluster-spins $\eta_{1},\eta_{2},\ldots,\eta_{\ell}$ of inner clusters are conditionally independent with the common distribution $\mu_{FKSW}(\eta_{j}=1|e)=\mu_{FKSW}(\eta_{j}=0|e)=1/2$. In summary, the spin $\sigma_{i}$ at the vertex $i$ is fixed according to $S(i)$ and its cluster-spin $\eta_{k}$. Given the edge configuration $e$, the function $S$ is fixed, the cluster-spins of inner clusters form Rademacher series and the cluster-spins of boundary clusters are determined by boundary conditions.
3 General lower bounds
In this section, we obtain an lower bound for $Q_{n}$ which holds for any random vector taking values in $\{-1,1\}^{n}$.
Theorem 3.1.
Consider an arbitrary random vector $\varepsilon=(\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n})$ taking values in $\{-1,1\}^{n}$. As before, we define
$Q_{n}=\sup_{x\in\mathbb{R}}\sup_{|v_{1}|,|v_{2}|,\ldots,|v_{n}|\geq 1}P(\varepsilon_{1}v_{1}+\varepsilon_{2}v_{2}+\cdots+\varepsilon_{n}v_{n}\in(x-1,x+1)).$
Then, for all $n\geq 1$, we have that
$Q_{n}\geq\binom{n}{[n/2]}2^{-n}.$ (9)
Proof.
Let $\tau_{1},\tau_{2},\ldots,\tau_{n}$ be independent Rademacher random variables such that
$P(\tau_{i}=1)=P(\tau_{i}=-1)=1/2,\forall i=1,2,\ldots,n.$
Write $\tau=(\tau_{1},\tau_{2},\ldots,\tau_{n})$. By definition of $Q_{n}$, we have that
$Q_{n}\geq\sup_{x\in\mathbb{R}}P(\sum_{i=1}^{n}\varepsilon_{i}\tau_{i}\in(x-1,x+1)|\tau).$
Hence, we have that
$Q_{n}\geq\sup_{x\in\mathbb{R}}P(\sum_{i=1}^{n}\varepsilon_{i}\tau_{i}\in(x-1,x+1)).$
Conditionally on the value of $\varepsilon$ in $\{-1,1\}^{n}$, the conditional distribution of the random vector $(\varepsilon_{1}\tau_{1},\varepsilon_{2}\tau_{2},\ldots,\varepsilon_{n}\tau_{n})$ is the same as the unconditional distribution of the Rademacher series $(\tau_{1},\tau_{2},\ldots,\tau_{n})$. Hence, the distribution of $(\varepsilon_{1}\tau_{1},\varepsilon_{2}\tau_{2},\ldots,\varepsilon_{n}\tau_{n})$ is the same as the Rademacher series $(\tau_{1},\tau_{2},\ldots,\tau_{n})$. Therefore, $\forall x\in\mathbb{R}$, we have that
$P(\sum_{i=1}^{n}\varepsilon_{i}\tau_{i}\in(x-1,x+1))=P(\sum_{i=1}^{n}\tau_{i}\in(x-1,x+1)).$
Then, it is not hard to see that
$Q_{n}\geq\sup_{x\in\mathbb{R}}P(\sum_{i=1}^{n}\tau_{i}\in(x-1,x+1))=\binom{n}{[n/2]}2^{-n}.\qed$
4 Upper bounds
As explained at the beginning of Section 2, without loss of generality, we assume that the external field $h$ vanishes. Recall Edwards-Sokal coupling and the notation in Section 2.
Our start point is the following observation:
Lemma 4.1.
The random variable $\sum_{i=1}^{n}\varepsilon_{i}v_{i}$ equals to the random variable $\sum_{k=1}^{\ell}\eta_{k}w_{k}+a$, where for $k=1,2,\ldots,\ell+q$, $w_{k}:=\sum_{j\in C_{k}}S(j)v_{j}$ and $a:=\sum_{k=\ell+1}^{\ell+q}\eta_{k}w_{k}=\sum_{k=\ell+1}^{\ell+q}\sum_{j\in C_{k}}\varepsilon_{j}v_{j}$. In particular,
$P\left(\sum_{i=1}^{n}\varepsilon_{i}v_{i}\in(x-1,x+1)\right)=\mu_{FKSW}\left(\sum_{k=1}^{\ell}\eta_{k}w_{k}=(x-a-1,x-a+1)\right).$
Note that the number $\ell$ of inner clusters and the real numbers $w_{1},w_{2},\ldots,w_{\ell}$ and $a$ are determined by the edge configuration $e$. Given the edge configuration $e$, the cluster-spins $\eta_{1},\eta_{2},\ldots,\eta_{\ell}$ are i.i.d. Bernoulli random variables with parameter $1/2$. By the classical Littlewood-Offord theorem [Erd45, Theorem 1] for Rademacher series, if $|w_{1}|,|w_{2}|,\ldots,|w_{\ell}|\geq 1$, then there exists a finite universal constant $C_{1}$ such that for all $x\in\mathbb{R}$,
$\mu_{FKSW}\left(\left.\sum_{k=1}^{\ell}\eta_{k}w_{k}\in(x-a-1,x-a+1)\right|e\right)\leq C_{1}/\sqrt{\ell}.$
However, although $|v_{1}|,|v_{2}|,\ldots,|v_{n}|\geq 1$, we don’t necessarily have
$|w_{k}|\geq 1,\forall k=1,2,\ldots,\ell$
in general. Fortunately, if the inner cluster $C_{k}$ consists of a single vertex, we do have $|w_{k}|\geq 1$ as $|w_{k}|$ equals to some $|v_{j}|$. By reordering the clusters if necessary, without loss of generality, we assume that $C_{1},C_{2},\ldots,C_{N}$ are all the inner clusters formed by a single vertex, where $N\leq\ell$ is the number of isolated (inner) vertices in the graph with the vertex set $V_{n}$ and the edge configuration $e$. By the classical Littlewood-Offord theorem for Rademacher series, there exists a finite universal constant $C_{1}$ such that for all $y\in\mathbb{R}$,
$\mu_{FKSW}\left(\left.\sum_{k=1}^{N}\eta_{k}w_{k}\in(y-1,y+1)\right|e\right)\leq C_{1}/\sqrt{N+1}.$
Conditionally on the edge configuration $e$, $(\eta_{1},\eta_{2},\ldots,\eta_{N})$ is independent from $(\eta_{N+1},\eta_{N+2},\ldots,\eta_{\ell})$. Hence, we have that
$\mu_{FKSW}\left(\left.\sum_{k=1}^{\ell}\eta_{k}w_{k}\in(x-a-1,x-a+1)\right|e\right)\\ =E\left(\mu_{FKSW}\left(\left.\left.\sum_{k=1}^{N}\eta_{k}w_{k}\in(y-1,y+1)\right|e,\eta_{N+1},\ldots,\eta_{\ell}\right)\right|e\right)\\ \leq E(C_{1}/\sqrt{N+1}|e)\leq C_{1}/\sqrt{N+1},$ (10)
where $y=x-a-\sum_{k>N}\eta_{k}w_{k}$. Typically, $N$ is of the same order as $n$. More precisely, we have the following result.
Lemma 4.2.
Recall (4). Then, there exist universal constants $c_{1}=c(K)$ and $C_{2}=C_{2}(K)$ such that for all $n$,
$P(N\geq c_{1}n)\geq 1-C_{2}n^{-10}.$
Combining previous results, we have that
$Q_{n}\leq C_{2}n^{-10}+C_{1}/\sqrt{c_{1}n}\leq C_{3}/\sqrt{n},$
where $C_{3}=C_{3}(K)<\infty$ is a universal constant.
Finally, it remains to prove Lemma 4.2.
Proof of Lemma 4.2.
We first sample the spin configuration $\sigma$ according to the Gibbs measure $\mu_{n}$ of the Ising model. Then, we sample the edge configuration $e$ as follows: For an edge $\{i,j\}\in\mathcal{E}_{n}^{+}$, if $\sigma_{i}\neq\sigma_{j}$, then set $e_{ij}=0$; if $\sigma_{i}=\sigma_{j}$, then set $e_{ij}=0$ with conditional probability $1-p_{ij}$, conditionally independent from the other edges. For an edge $\{i,j\}\in\mathcal{E}_{n}^{-}$, if $\sigma_{i}=\sigma_{j}$, then set $e_{ij}=0$; if $\sigma_{i}\neq\sigma_{j}$, then set $e_{ij}=0$ with conditional probability $1-p_{ij}$, conditionally independent from the other edges. By Edwards-Sokal coupling explained in Section 2, the joint distribution of $(\sigma,e)$ is precisely $\mu_{FKSW}$. Define
$V_{n}^{+}=\{i\in V_{n}:\sigma_{i}=1\},\quad V_{n}^{-}=\{i\in V_{n}:\sigma_{i}=-1\}.$
Then, $\{V_{n}^{+},V_{n}^{-}\}$ is a partition of $V_{n}$ and $\max(|V_{n}^{+}|,|V_{n}^{-}|)\geq n/2$. Without loss of generality, we may assume $|V_{n}^{+}|\geq n/2$. Consider the conditioned bond percolation process, let $N^{+}$ be the number of isolated vertices in $V_{n}^{+}$. It suffices to prove Lemma 4.2 for $N^{+}$ instead of $N$. For this purpose, we consider the following percolation process on the graph $G=(W,\mathcal{E})$. Here, $W=\overline{V_{n}}$ and
=$\mathcal{E}=\{\{i,j\}\in\mathcal{E}_{n}^{+}:\sigma_{i}=\sigma_{j}\}\cup\{\{i,j\}\in\mathcal{E}_{n}^{-}:\sigma_{i}\neq\sigma_{j}\}.$
Let $V$ be a subset of vertices of $W$. Assume that $|V|=m\geq n/2$. Here, $V=V_{n}^{+}$. For the probability $p_{ij}$ that $\{i,j\}\in\mathcal{E}$ is open, we assume that there exists a universal constant $K<\infty$ such that for all $i\in V$,
$\sum_{j:\{i,j\}\in\mathcal{E}}-\ln(1-p_{ij})\leq K,$ (11)
which is guaranteed by (4) and $p_{ij}=1-e^{-|J_{ij}|}$. Let $M$ be the number of isolated vertices in $V$. We have the following claim:
Claim 4.3.
There exist universal constants $c=c(K)$ and $C=C(K)$ such that
$P(M>cm)\geq 1-Cm^{-10}.$ (12)
To prove (12), we gradually discover all the isolated vertices in $V$ by the following algorithm. We list the vertex set $V$ as $i_{1},i_{2},\ldots,i_{m}$. Let $V_{1}=V$ and $v_{1}=i_{1}$. Next, we sample the edges adjacent to $i_{1}$ according to Bernoulli random variables with parameter $p_{i_{1}j}$, i.e. we keep the edge $\{i_{1},j\}$ with probability $p_{i_{1}j}$ and remove the edge $\{i_{1},j\}$ with probability $1-p_{i_{1}j}$. And there is independence among different edges. Define
$U_{1}=\{j\in W:\{i_{1},j\}\text{ is not removed}\}.$
Note that $i_{1}$ is isolated iff $U_{1}=\emptyset$. If $U_{1}\neq\emptyset$, then we immediately know that the vertices in $U_{1}\cup\{i_{1}\}$ are not isolated. In any case, we don’t need to examine the status of the vertices in $U_{1}\cup\{i_{1}\}$ anymore in the following steps. We will recursively define $V_{k}$, $v_{k}\in V_{k}$ and $U_{k}$: The set $V_{k}$ is the unexplored region of $V$ after the first $k-1$ steps. In the $k$-th step, we choose some vertex $v_{k}\in V_{k}$ and examine the remainder edges adjacent to $v_{k}$. Roughly speaking, the vertices in $U_{k}$ are non-isolated vertices discovered in the $k$-th step. Suppose $V_{1},V_{2},\ldots,V_{k},v_{1},\ldots,v_{k},U_{1},U_{2},\ldots,U_{k}$ are well defined. We set $V_{k+1}=V_{k}\setminus(U_{k}\cup\{v_{k}\})$. If $V_{k+1}=\emptyset$, then the algorithm actually finishes and we set $U_{k+1}=V_{k+2}=U_{k+2}=V_{k+3}=\cdots=\emptyset$. If $V_{k+1}\neq\emptyset$, we choose the vertex $i_{f(k+1)}\in V_{k+1}$ with the least index $f(k+1)$ and define $v_{k+1}=i_{f(k+1)}$. Then, we sample the remainder edges adjacent to $v_{k+1}$. We remove an edge $e$ with probability $1-p_{e}$. (Note that the edges $\{v_{j},v_{k+1}\}$ are already removed for $j=1,2,\ldots,k$. Otherwise, $v_{k+1}$ will belong to some $U_{j}$ and it will not belong to the unexplored region $V_{k+1}$.) Define
$U_{k+1}=\{j\in W:\{v_{k+1},j\}\text{ is not removed}\}.$
Note that $v_{k+1}$ is isolated iff $U_{k+1}=\emptyset$. Let
$T=\max\{k\geq 1:V_{k}\neq\emptyset\}.$
We say that the algorithm lasts for $T$ turns. By the construction of the algorithm, we have the following expression for the number $M$ of isolated vertices:
$M=\sum_{k=1}^{T}1_{U_{k}=\emptyset}.$
Since we may discover the set of isolated vertices by this algorithm, we will call such an algorithm the algorithm to search (all) the isolated vertices. Define
$X_{k}=|U_{k}|,\mathcal{F}_{k}=\sigma(V_{j},v_{j},U_{j},j\leq k)\text{ for }k\geq 1.$
Define $\mathcal{F}_{0}=\{\emptyset,\Omega\}$ be the trivial minimal $\sigma$-field. Then, we have that
$\sum_{k=1}^{T}(X_{k}+1)\geq m=|V|.$
We have an inequality here for two reasons: Firstly, $U_{k}$ may contain vertices in $W\setminus V$; secondly, there may exist overlaps between different $U_{k}$. For a Bernoulli random variable $\xi$ such that $P(\xi=1)=1-P(\xi=0)=p$, $\xi$ is stochastically dominated by a Poisson random variable with parameter $\lambda=-\ln(1-p)$. Besides, the sum of independent Poisson random variables is still a Poisson random variable with the new parameter as the sum of parameters. Hence, conditionally on $\mathcal{F}_{k}$, $X_{k}$ is stochastically dominated by a Poisson random variable with parameter $K$, where $K$ appears in the condition (11). Let $(Y_{k})_{k\geq 1}$ be i.i.d. Poisson random variables with the same parameter $K$. Note that $T$ stochastically dominates $T^{Y}$, where
$T^{Y}=\inf\{k\geq 1:Y_{1}+Y_{2}+\cdots+Y_{k}\geq m-k\}.$
A key observation is that the number $M$ of isolated vertices stochastically dominates
$M^{Y}=\sum_{k=1}^{T^{Y}}1_{Y_{k}=0}.$
By standard arguments with exponential Markov inequality for Poisson random variables, there exist universal constants $c=c(K)$ and $C=C(K)$ such that
$P(T^{Y}\leq cm)\leq Ce^{-m}\text{ and }P(M^{Y}\geq cm)\geq 1-Cm^{-10}.$
Finally, by the key observation,
$P(M\geq cm)\overset{M^{Y}\leq_{s.t.}M}{\geq}P(M^{Y}\geq cm)\geq 1-Cm^{-10}$
and (12) is proved. ∎